Tyrion:

The aquarium has 6 fewer yellow fish than green fish. 40 percent of the fish are yellow. How many green fish are in the aquarium? Show your work.

dizzymustang156:

answer choices?

Tyrion:

ion need the answer i need the work

HaydenwH:

now dizzy work him through it instead of doin it for him this time ?

lilcutie2:

what's 40 minus 6

Tyrion:

34

dizzymustang156:

wow just wow

Tyrion:

so no one kan help??

dizzymustang156:

just go with the answer I told you please

lilcutie2:

T=Total G=Green Y=Yellow T=G+Y (given) G=Y+6 (given) T=(Y+6)+Y (substitute for G) T=2Y+6 Y=.4*T (given) T=2*(.4*T) + 6 (substitute for Y) Solve for T T=.8T + 6 .2T=6 T=30 Solve for G G=.6*T (given) G=18

dizzymustang156:

I JUST SHOWED HIM Y DID SOMEONE DELETE IT

HaydenwH:

alright ty. let the number of green fish =G the number of yellow fish =G-6 0.4 is the decimal of 40% the equation: 0.4[g+g-6]=g-6, now solve for g

HaydenwH:

i posted wht got deleted, i deleted it bc i was fixing the equation, but u shouldnt have gave him the answer, let him work it

dizzymustang156:

ugh

HaydenwH:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @HaydenwH alright ty. let the number of green fish =G the number of yellow fish =G-6 0.4 is the decimal of 40% the equation: 0.4[g+g-6]=g-6, now solve for g \(\color{#0cbb34}{\text{End of Quote}}\) @Tyrion

dizzymustang156:

Im never answering questions again people will just delete the whoole THING!!!!!!

Tyrion:

i got it already

HaydenwH:

igh

Tyrion:

the question is different now